OpenStudy (anonymous):

graph the equation by plotting 3 points -y=x+1

OpenStudy (anonymous):

-y = x + 1 If x = 0. - y = 0 + 1 = 1 y = -1 (0,-1) If y = 0, - 0 = x + 1, 0 = x + 1, x = - 1. (-1,0). If x = 1, - y = 1 + 1 = 2, y = -2. (1,-2).

OpenStudy (anonymous):

thanks can you help me again

OpenStudy (anonymous):

2y=-x+2

OpenStudy (anonymous):

x = 0 2y = -0 +2 2y=2 y=1 (0,1) y=0 2*0=-x+2 0=-x+2 x=2 (2,0) x=1 2y=-1+2 2y=1 y=1/2 (1,1/2)